# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

### v1

@@ II-C Speech Vocoders
Vocoders are essential for converting acoustic features such as mel-spectrograms into intelligible audio waveforms and are vital in determining the naturalness and quality of synthesized speech. We broadly categorize existing vocoders according to their model architectures, i.e., RNN-, CNN-, GAN-, and diffusion-based vocoders.
@@
CNN-based vocoders. By leveraging the parallel nature of convolutional operations, CNN-based vocoders [73, 140, 141] can generate high-quality speech more efficiently, making them ideal for real-time applications. A key strength of CNN-based vocoders is their ability to balance synthesis quality and efficiency. However, they often require extensive training data and careful hyperparameter tuning to achieve optimal performance. WaveNet [73] is a probabilistic autoregressive model that generates waveforms sample by sample conditioned on all preceding samples and auxiliary inputs, such as linguistic features and mel-spectrograms. It employs stacks of dilated causal convolutions, enabling long-range dependence modeling in speech signals without relying on recurrent connections. Parallel WaveNet [140] addresses WaveNet’s inference speed limitations while maintaining comparable synthesis quality. It introduces a non-autoregressive mechanism based on a teacher-student framework, where the original WaveNet (teacher) distills knowledge into a student model. The student generates samples in parallel, enabling real-time synthesis without waveform quality degradation.
@@
GAN-based vocoders. GANs have been widely adopted in vocoders for high-quality speech generation [142, 143, 116, 144, 145], leveraging adversarial losses to improve realism. GAN-based vocoders typically consist of a generator that produces waveforms conditioned on acoustic features, such as mel-spectrograms, and a discriminator that distinguishes between real and synthesized waveforms. Models like Parallel WaveGAN [144] and HiFi-GAN [116] have demonstrated the effectiveness of GANs in vocoding by introducing tailored loss functions, such as multi-scale and multi-resolution spectrogram losses, to ensure naturalness in both time and frequency domains. These models can efficiently handle the complex, non-linear relationships inherent in speech signals, resulting in high-quality synthesis. A key advantage of GAN-based vocoders is their parallel inference capability, enabling real-time synthesis with lower computational costs compared to autoregressive models. However, training GANs can be challenging due to instability and mode collapse. Despite these challenges, GAN-based vocoders continue to advance the state-of-the-art in neural vocoding, offering a compelling combination of speed and audio quality.
@@
Diffusion-based vocoders. Inspired by diffusion probabilistic models [146] that have shown success in visual generation tasks, diffusion-based vocoders [147, 113, 148, 149, 150] present a novel approach to natural-sounding speech synthesis. The core mechanism of diffusion-based vocoders involves two stages: a forward process and a reverse process. In the forward process, clean speech waveforms are progressively corrupted by adding noise in a controlled manner, creating a sequence of intermediate noisy representations. During training, the model learns to reverse this process, progressively denoising the corrupted signal to reconstruct the original waveform. Diffusion-based vocoders, such as WaveGrad [149] and DiffWave [148], have demonstrated remarkable performance in generating high-fidelity waveforms while maintaining temporal coherence and natural prosody. They offer advantages over previous vocoders, including robustness to over-smoothing [151] and the ability to model complex data distributions. However, their iterative sampling process can be computationally intensive, posing challenges for real-time applications.
@@
Other vocoders. There are also many other types of vocoders such as flow-based [152, 153, 154, 155, 156] and VAE-based vocoders [157, 158, 159]. These methods provide unique strengths for speech synthesis such as efficiency and greater flexibility in modeling complex speech variations. Readers can refer to the survey paper from Tan et al. [42] for more details.
@@ II-D Fully End-to-end TTS models
Fully end-to-end TTS methods [160, 161, 76, 159, 162] directly generate speech waveforms from textual input, simplifying the “acoustic model → vocoder” pipeline and achieving efficient speech generation. Char2Wav [160] is an early neural text-to-speech (TTS) system that directly synthesizes speech waveforms from character-level text input. It integrates two components and jointly trains them: a recurrent sequence-to-sequence model with attention, which predicts acoustic features (e.g., mel-spectrograms) from text, and a SampleRNN-based neural vocoder [136] that generates waveforms from these features. Similarly, FastSpeech 2s [76] directly synthesizes speech waveforms from texts by extending FastSpeech 2 [76] with a waveform decoder, achieving high-quality and low-latency synthesis.   VITS [159] is another fully end-to-end TTS framework. It integrates a variational autoencoder (VAE) with normalizing flows [163] and adversarial training, enabling the model to learn latent representations that capture the intricate variations in speech, such as prosody and style. VITS combines non-autoregressive synthesis with stochastic latent variable modeling, achieving real-time waveform generation without compromising naturalness. There are more end-to-end TTS models such as Tacotron [74], ClariNet [161], and EATS [162], refer to another survey [42] for more details. End-to-end controllable methods that emerged in recent years will be discussed in Section IV.
TABLE II: Popular open-source speech quantization methods.
SSCP: Self-supervised context prediction, RVQ: Residual vector quantization [168].
@@ II-E Acoustic Feature Representations
Continuous representations. Continuous representations (e.g., mel-spectrograms) of intermediate acoustic features use a continuous feature space to represent speech signals. These representations often involve acoustic features that capture frequency, pitch, and other characteristics without discretizing the signal. The advantages of continuous features are: 1) Continuous representations retain fine-grained detail, enabling more expressive and natural-sounding speech synthesis. 2) Since continuous features inherently capture variations in tone, pitch, and emphasis, they are well-suited for prosody control and emotional TTS. 3) Continuous representations are more robust to information loss and can avoid quantization artifacts, allowing smoother, less distorted audio. GAN-based [116, 144, 145] and diffusion-based methods [148, 147] often utilize continuous feature representations, i.e., mel-spectrograms. However, continuous representations are typically more computationally demanding and require larger models and memory, especially in high-resolution audio synthesis.
Discrete tokens. In discrete token-based TTS, the intermediate acoustic features (e.g., quantized units or phoneme-like tokens) are discrete values, similar to words or phonemes in languages. These are often produced using quantization techniques or learned embeddings, such as HuBERT [166] and SoundStream [168]. The advantages of discrete tokens are: 1) Discrete tokens can encode phonemes or sub-word units, making them concise and less computationally demanding to handle. 2) Discrete tokens often allow TTS systems to require fewer samples to learn and generalize, as the representations are compact and simplified. 3) Using discrete tokens simplifies cross-modal TTS applications like voice cloning or translation-based TTS, as they map well to text-like representations such as LLM tokens. LLM-based [78, 103, 106, 105] and zero-shot TTS methods [17, 78, 87] often adopt discrete tokens as their acoustic features. However, discrete representation learning may result in information loss or lack the nuanced details that can be captured in continuous representations.
Table IV and III summarize the types of acoustic features of representative methods. Table II summarizes popular open-source speech quantization methods.
@@ III Uncontrollable TTS
The development of Uncontrollable Text-To-Speech (UC-TTS) systems represents a significant shift from traditional, linguistics-based synthesis to modern, data-driven deep learning techniques. This shift highlights the integration of both local and global information to produce speech with human-like quality and naturalness. This survey explores UC-TTS evolution, emphasizing the role of local and global information in enhancing speech fidelity and expressiveness.
In the context of UC-TTS, ”uncontrollable” refers to the absence of explicit control mechanisms for speech features such as emotion, timbre, and speaker style. Despite this lack of explicit control, the goal is to achieve natural, fluid speech while minimizing issues like mispronunciations and omissions.
@@ III-A Early Approaches: Statistical Models
Early Text-To-Speech (TTS) systems relied on statistical models such as Hidden Markov Models (HMMs) [64, 65] and early neural network-based parametric methods [111, 112]. These models operated at the frame level, using acoustic models and vocoders for text-to-speech conversion. Notable contributions from Tokuda et al. [173] employed HMMs for statistical parametric synthesis, focusing on local features like phonemes, accents, and prosody to improve speech naturalness.
@@ III-B Sequence-to-Sequence Models
The emergence of sequence-to-sequence models represents a significant breakthrough by removing the need for explicit linguistic features, thereby enabling the capture of the nuances and idiosyncrasies of human speech. Models such as Tacotron [74] and Tacotron 2  [175] utilize recurrent neural networks (RNNs) with attention mechanisms to effectively model the complex, nonlinear nature of speech sequences. These innovations allow for precise tuning of speech parameters, enhancing prosody and rhythm by modeling entire utterances rather than isolated phonetic units.
Building on these advancements, Deep Voice 3[176] introduces a fully convolutional sequence-to-sequence architecture that significantly accelerates training speed compared to RNN-based models. This approach achieves training times an order of magnitude faster, enabling scalability to handle large datasets. Additionally, the use of a position-augmented attention mechanism in Deep Voice 3 enhances the naturalness of synthesized speech, achieving competitive mean opinion scores, especially when paired with advanced neural vocoders like WaveNet. This development not only improves training efficiency but also enhances the scalability and naturalness of text-to-speech systems.
@@ III-C Transformer-Based Models
III-D Advanced Architectures: Integrating Flow and Diffusion Models
Recent advancements have shifted towards integrating global information within end-to-end architectures to enhance speech naturalness and coherence. Flow-based models like Glow-TTS[133] and Flow-TTS[132] exemplify this by employing invertible transformations that maintain the balance between local precision and global coherence. These architectures enable the synthesis of high-fidelity speech by modeling complex dependencies across the entire utterance, thus improving the overall fluidity and naturalness of the generated speech.
Moreover, the introduction of diffusion models in TTS, such as WaveGrad 2[178], highlights the shift towards models that can iteratively refine speech output. These models use score matching and diffusion processes to generate speech directly from phoneme sequences, effectively capturing both local nuances and overarching global patterns. The iterative nature of these models allows for adjustments that enhance the quality of the synthesized audio, accommodating variations in speech without explicit control over specific attributes.
The integration of adversarial training and variational autoencoders (VAEs) further exemplifies the evolution towards incorporating global information. Systems like VITS[159] leverage these techniques to enhance expressiveness and naturalness by learning complex mappings between text and speech. This approach allows the model to manage variations in prosody and rhythm inherently derived from the textual input, aligning with the objectives of UC-TTS to produce diverse and natural speech outputs.
@@ IV Controllable TTS
In this section, we first review recent TTS work from the perspective of model architecture, followed by a detailed discussion of control modes in controllable TTS.
IV-A Model Architectures
@@
Current model architectures can be broadly classified into two main categories: the first is the non-autoregressive (NAR) generative models, which are based on HMMs, neural networks, VAEs, diffusion models, flow matching, and other NAR techniques. The second category relies on autoregressive (AR) codec language models, which typically quantize speech into discrete tokens and use decoder-only models to autoregressively generate these tokens. We summarize the NAR-based and AR-based controllable TTS methods in Table III and Table IV, respectively.
IV-A1 Fully Non-Autoregressive (NAR) Architectures
HMM-based Approaches.  In the realm of Controllable Text-To-Speech (CTTS), advancements in Hidden Markov Model (HMM) architectures have significantly enhanced the manipulation of speech elements such as emotion and prosody. Yamagishi et al. [70] pioneered this field by introducing style-dependent and style-mixed modeling, which allowed precise emulation of human-like emotional nuances and versatile synthesis across various styles by incorporating style as a contextual variable. Building on this foundation, Qin et al.[179] developed the ”average emotion model,” which utilized MLLR-based adaptation to modulate emotions like happiness and sadness even with limited data, thus advancing the emotional intelligence of synthetic speech systems.
Furthering expressive variability, Nose et al.[119] integrated subjective style intensities and a multiple-regression global variance model into HMM frameworks, addressing over-smoothing and enabling nuanced emotional expressions. Lorenzo-Trueba et al.[72] expanded on these capabilities with CSMAPLR adaptation, introducing ”emotion transplantation” to transfer emotional states between speakers while preserving voice distinctiveness, enhancing personalized human-computer interaction. These innovations in HMM architectures have broadened the expressiveness and individuality in synthetic speech, augmenting technological interfaces and paving the way for future developments in adaptive, lifelike speech solutions.
Transformer-based Approaches.  Advancements in Controllable Text-to-Speech (TTS) technology highlight the integration of deep learning with audio processing, driven by Transformer-based architectures. Ren et al.[15] introduced FastSpeech, a feed-forward non-autoregressive Transformer model that significantly enhances TTS efficiency by reducing inference time and improving the stability issues found in autoregressive models like Tacotron 2. This model provides precise control over prosodic features through duration prediction, effectively tackling the one-to-many mapping challenge. FastSpeech 2[180] builds on this by integrating pitch and energy control, eliminating the need for the complex teacher-student distillation process, thus enhancing training efficiency and improving voice quality. Parallel Tacotron[84] further advances TTS by employing a variational autoencoder-based residual encoder, capturing intricate prosodic nuances. This approach, combined with iterative spectrogram loss, significantly enhances the naturalness and quality of synthesized speech. Additionally, FastPitch[77] incorporates direct pitch prediction into its architecture, enabling fully parallelized synthesis and precise pitch manipulation. This capability enhances expressiveness and retains the efficiency benefits established by FastSpeech. These innovations significantly contribute to the development of more interactive and natural AI-driven communication systems, underscoring the potential of integrating AI with human-centric disciplines to craft a future where technology and humanity coexist harmoniously.
VAE-based Approaches.  Recent advancements in Controllable Text-To-Speech (TTS) systems are largely driven by the integration of Variational Autoencoder (VAE) architectures, which enhance the flexibility and precision of speech modulation. Zhang et al.[18] pioneered the use of VAEs in end-to-end speech synthesis, creating disentangled latent representations that allow effective style control and transfer, especially in prosody and emotion management, outperforming the Global Style Token model in style transfer tasks. Building on this, Hsu et al.[130] developed a hierarchical generative model with a conditional VAE framework and a Gaussian mixture model, enabling precise control over complex speech attributes such as environment and style, thus improving expressive speech synthesis through refined noise and speaker characteristic management. Liu et al.[181] further advanced the field with the CLONE model, a single-stage TTS system that resolves the one-to-many mapping issue and enhances high-frequency information reconstruction. By employing a conditional VAE with normalizing flows and a dual path adversarial training mechanism with multi-band discriminators, CLONE achieves nuanced control over prosody and energy, demonstrating superior performance in both speech quality and prosody control compared to state-of-the-art models. These collective innovations highlight the adaptability of VAEs in managing complex speech generation tasks, marking significant progress toward more dynamic and versatile TTS technologies, with ongoing research promising even greater advancements.
Diffusion-based Approaches. The core concept of diffusion-based models is to generate target data by progressively removing noise. During the forward diffusion phase, noise is incrementally added to the original data to form a noise distribution. In the generation phase, a reverse denoising process is employed to gradually recover high-quality speech from the noise. Grad-StyleSpeech [182] introduces a hierarchical transformer encoder to create a representative noise prior distribution for speaker-adaptive settings using score-based diffusion models. NaturalSpeech 2 [86] uses a neural audio codec with residual vector quantizers to obtain quantized latent vectors, which are then generated using a diffusion model conditioned on text input. NaturalSpeech 3 [87] decomposes speech into distinct subspaces that represent different attributes and generates each subspace independently. DEX-TTS [183] improves DiT-based diffusion networks by applying overlapping patchify and convolution-frequency patch embedding strategies. E3 TTS [184] models the temporal structure of the waveform through the diffusion process, eliminating the need for any intermediate representations such as spectrogram features or alignment information.
Applying diffusion models to TTS requires a complex pipeline due to the need for precise temporal alignment between text and speech and the high fidelity required for audio data. This includes domain-specific modeling, such as phoneme and duration [86]. To address the issue of reduced naturalness caused by the addition of duration models, DiTTo-TTS [185] leverages the off-the-shelf pre-trained text and speech encoders without relying on speech domain-specific modeling by incorporating cross-attention mechanisms with the prediction of the total length of speech representations. Similarly, SimpleSpeech [186] proposes a speech codec model (SQ-Codec) based on scalar quantization and uses the sentence duration to control the generated speech length.
Flow-based Approaches. Flow-based methods leverage invertible flow transformations to learn mappings from target speech features to simple distributions, typically standard Gaussian distributions. Due to their invertibility, this mechanism can directly sample from the simple distribution and generate high-fidelity speech in the reverse direction. Audiobox [187] and P-flow [188] employ non-autoregressive flow-matching models for efficient and stable speech synthesis. VoiceBox [189] also employs a flow-matching to generate speech, effectively casting the TTS task into a speech infilling task. SpeechFlow [190] is trained on 60k hours of untranscribed speech with flow matching and mask conditions and can be fine-tuned with task-specific data to match or surpass existing expert models. This highlights the potential of generative models as foundation models for speech applications. HierSpeech++ [191] proposes a hierarchical variational inference method. FlashSpeech [192] is built on a latent consistency model and applies a novel adversarial consistency training approach that can train from scratch without the need for a pre-trained diffusion model as the teacher, achieving speech generation in one or two steps.
Recently, E2 TTS [193] converts text input into a character sequence with filler tokens and trains a mel spectrogram generator based on audio infilling task, achieving human-level naturalness. Inspired by E2 TTS, F5-TTS [194] refines the text representation with ConvNext v2 [195], facilitating easier alignment with speech. E1 TTS [196] further distills a diffusion-based TTS model into a one-step generator with distribution matching distillation [197, 198], reducing the number of network evaluations in sampling from diffusion models. SimpleSpeech 2 [199] introduces a flow-based scalar transformer diffusion model. The work also provides a theoretical analysis, showing that the inclusion of a small number of noisy labels in a large-scale dataset is equivalent to introducing classifier-free guidance during model optimization.
Other NAR Approaches. Other works leverage GAN-based or Masked Generative-based methods for TTS generation. StyleTTS 2 [89] employs large pre-trained speech language models (SLMs) such as Wav2Vec 2.0 [165], HuBERT [166], and WavLM [200] as discriminators, in combination with a novel differentiable duration modeling approach. This setup uses SLM representations to enhance the naturalness of the synthesized speech. MaskGCT [78] proposes masked generative transformers without requiring text-speech alignment supervision and phone-level duration. The model employs a two-stage system, both trained using a mask-and-predict learning paradigm.
IV-A2 Autoregressive (AR) Architectures
RNN-based Approaches. Controllable Text-To-Speech (TTS) technology has seen significant advancements through innovations in neural network architectures, facilitating speech that is both natural-sounding and adaptable in terms of emotion, prosody, and pitch. A key breakthrough was the introduction of Tacotron[74], a sequence-to-sequence architecture that effectively integrates prosodic variations, laying the groundwork for precise control over speech attributes. Tacotron 2[175] further enhanced this capability by better managing prosodic variability, though it averaged these variations, indicating a need for more sophisticated control methods. To address these constraints, Wang et al.[19] introduced Global Style Tokens (GSTs), using an unsupervised approach to encapsulate diverse speech styles into fixed tokens, thus enabling versatile style transfer within the Tacotron framework. Skerry-Ryan et al.[201] further advanced this by incorporating prosodic embeddings, providing detailed control over timing and intonation, and significantly improving the replication of emotional expressions in synthetic speech.
LLM-based Approaches. Inspired by the success of large language models (LLMs) in natural language processing (NLP), recent studies have explored leveraging in-context learning for zero-shot TTS generation.
Building on VALL-E, various improvements have been proposed. VALL-E X [207] extends VALL-E to multilingual scenarios, supporting zero-shot cross-lingual speech synthesis and speech-to-speech translation. ELLA-V [208] introduces a sequence order rearrangement step, enhancing local alignment between phoneme and acoustic modalities. RALL-E [209] incorporates prosody tokens as chain-of-thought prompting [210] to stabilize the generation of speech tokens. VALL-E R [211] improves phoneme-to-acoustic alignment and adopts codec-merging to boost decoding efficiency and reduce computational overhead. VALL-E 2 [212] introduces repetition aware sampling and grouped code modeling for greater stability and faster inference. HALL-E [213] adopts a hierarchical post-training framework, effectively managing the trade-off between reducing frame rate and producing high-quality speech.
Beyond the foundational improvements introduced by VALL-E and its immediate extensions, further advancements have focused on enhancing speech alignment, quality, and robustness. SpearTTS [203] and Make-a-voice [204] use semantic tokens to bridge the gap between text and acoustic features. FireRedTTS [214] further optimizes the tokenizer architecture to enhance speech quality. CoFi-Speech [215] generates speech in a coarse-to-fine manner via a multi-scale speech coding and generation approach, producing natural and intelligible speech. CosyVoice [17] employs supervised semantic tokens to enhance content consistency and speaker similarity in zero-shot voice cloning. Similarly, BASE TTS [216] introduces discrete speech representations based on the WavLM [200] self-supervised model, focusing on phonemic and prosodic information. SeedTTS [217] also proposes a self-distillation method for speech decomposition and a reinforcement learning approach to enhance the robustness, speaker similarity and controllability of generated speech. Based on this framework, Bailing-TTS [218] enhances the alignment of text and speech tokens using a continual semi-supervised learning strategy, enabling high-quality synthesis of Chinese dialect speech.
Although models using discrete tokens as intermediate representations have achieved notable success in zero-shot TTS, they still face fidelity issues compared to the continuous representation like Mel spectrograms [219, 220]. MELLE [219] optimizes the training objectives and sampling strategy, marking the first exploration of using continuous-valued tokens instead of discrete-valued tokens within the paradigm of autoregressive speech synthesis models. Similar to MELLE, ARDiT [220] encodes audio as vector sequence in continuous space and autoregressively generates these sequences by a decoder-only transformer.
@@ IV-A2 Autoregressive (AR) Architectures
FastSpeech[15] Transformer WaveGlow MelS 2019.05
DWAPI [221] DNN Straight MelS + F0 + Intensity 2020.04
FastSpeech 2[180] Transformer Parallel WaveGAN MelS 2020.06
@@
Parallel Tacotron[222] Transformer + CNN WaveRNN MelS 2020.10
StyleTagging-TTS [223] Transformer + CNN HiFi-GAN MelS 2021.04
SC-GlowTTS [224] Transformer + Conv HiFi-GAN MelS 2021.06
Meta-StyleSpeech [225] Transformer MelGAN MelS 2021.06
DelightfulTTS[226] Transformer + CNN HiFiNet MelS 2021.11
YourTTS [82] Transformer HiFi-GAN LinS 2021.12
DiffGAN-TTS[227] Diffusion + GAN HiFi-GAN MelS 2022.01
StyleTTS [88] CNN + RNN + GAN HiFi-GAN MelS 2022.05
GenerSpeech [90] Transformer + Flow-based HiFi-GAN MelS 2022.05
NaturalSpeech 2 [86] Diffusion Codec Decoder Token 2022.05
Cauliflow[228] BERT + Flow UP WaveNet MelS 2022.06
CLONE[181] Transformer + CNN WaveNet MelS + LinS 2022.07
PromptTTS [101] Transformer HiFi-GAN MelS 2022.11

### v2

@@ II-C Speech Vocoders
Vocoders are essential for converting acoustic features, such as mel-spectrograms, into intelligible audio waveforms and are vital in determining the naturalness and quality of synthesized speech. We broadly categorize existing vocoders according to their model architectures, i.e., RNN-, CNN-, GAN-, and diffusion-based vocoders.
@@
CNN-based Vocoders. By leveraging the parallelism of convolutional operations, CNN-based vocoders [73, 139, 140] can achieve higher speech quality and more efficient synthesis compared to parametric vocoders [133, 134], making them ideal for applications that demand real-time and natural speech synthesis. However, they often require extensive training data and careful hyperparameter tuning to achieve optimal performance. WaveNet [73] is a probabilistic autoregressive model that generates waveforms sample by sample, conditioned on all preceding samples and auxiliary inputs, such as linguistic features and mel-spectrograms. It employs stacks of dilated causal convolutions, enabling long-range dependence modeling in speech signals without relying on recurrent connections. Parallel WaveNet [139] addresses WaveNet’s inference speed limitations while maintaining comparable synthesis quality. It introduces a non-autoregressive mechanism based on a teacher-student framework, where the original WaveNet (teacher) distills knowledge into a student model. The student generates samples in parallel, enabling real-time synthesis without waveform quality degradation.
@@
GAN-based Vocoders. GANs have been widely adopted in vocoders for high-quality speech generation [141, 142, 115, 143, 144], leveraging adversarial losses to improve realism. GAN-based vocoders typically consist of a generator that produces waveforms conditioned on acoustic features, such as mel-spectrograms, and a discriminator that distinguishes between real and synthesized waveforms. Models like Parallel WaveGAN [143] and HiFi-GAN [115] have demonstrated the effectiveness of GANs in vocoding by introducing tailored loss functions, such as multi-scale and multi-resolution spectrogram losses, to ensure naturalness in both time and frequency domains. These models can efficiently handle the complex, non-linear relationships inherent in speech signals, resulting in high-quality synthesis. A key advantage of GAN-based vocoders is their parallel inference capability, enabling real-time synthesis with lower computational costs compared to autoregressive models. However, training GANs can be challenging due to instability and mode collapse caused by imbalanced adversarial dynamics, vanishing gradients, and the generator overfitting to limited patterns [145, 146]. Despite these challenges, GAN-based vocoders continue to advance the SOTA in neural vocoding, offering a compelling combination of speed and audio quality.
@@
Diffusion-based Vocoders. Inspired by diffusion probabilistic models [147] that have shown success in visual generation tasks, diffusion-based vocoders [148, 112, 149, 150, 151] present an alternative approach to natural-sounding speech synthesis. The core mechanism of diffusion-based vocoders involves two stages: a forward process and a reverse process. In the forward process, clean speech waveforms are progressively corrupted by adding noise in a controlled manner, creating a sequence of intermediate noisy representations. During training, the model learns to reverse this process, progressively denoising the corrupted signal to reconstruct the original waveform. Diffusion-based vocoders, such as WaveGrad [150] and DiffWave [149], have demonstrated remarkable performance in generating high-fidelity waveforms while maintaining temporal coherence and natural prosody. They offer advantages over previous vocoders, including robustness to over-smoothing [152] and the ability to model complex data distributions. However, their iterative sampling process can be computationally intensive, posing challenges for real-time applications.
@@
Other Vocoders. There are also many other types of vocoders, such as flow-based [153, 154, 155, 156, 157] and VAE-based vocoders [122, 158, 159]. These methods provide unique strengths for speech synthesis, such as efficiency and greater flexibility in modeling complex speech variations. Readers can refer to the survey paper from Tan et al. [42] for more details.
@@ II-D Fully End-to-end TTS models
Fully end-to-end TTS methods [160, 161, 76, 159, 162] directly generate speech waveforms from textual input, simplifying the “acoustic model → vocoder” pipeline and achieving efficient speech generation. Char2Wav [160] is an early neural TTS system that directly synthesizes speech waveforms from character-level text input. It integrates two components and jointly trains them: a recurrent sequence-to-sequence model with attention, which predicts acoustic features (e.g., mel-spectrograms) from text, and a SampleRNN-based neural vocoder [135] that generates waveforms from these features. Similarly, FastSpeech 2s [76] directly synthesizes speech waveforms from texts by extending FastSpeech 2 [76] with a waveform decoder, achieving high-quality and low-latency synthesis. VITS [159] is another fully end-to-end TTS framework. It integrates a VAE with normalizing flows [163] and adversarial training, enabling the model to learn latent representations that capture the intricate variations in speech, such as prosody and style. VITS combines non-autoregressive synthesis with stochastic latent variable modeling, achieving real-time waveform generation without compromising naturalness. There are more end-to-end TTS models such as Tacotron [74], ClariNet [161], and EATS [162], readers can refer to another survey [42] for more details. End-to-end controllable methods that emerged in recent years will be discussed in Section IV.
TABLE II: Popular open-source speech quantization and tokenization methods.
SSCP: Self-supervised context (token) prediction, RVQ: Residual vector quantization [170].
@@ II-E Acoustic Feature Representations
Continuous Representations. Continuous representations (e.g., mel-spectrograms and VAE features) of intermediate acoustic features use a continuous feature space to model speech signals. These representations often involve acoustic features that capture frequency, pitch, and other characteristics without discretizing the signal. The advantages of continuous features are: 1) Continuous representations retain fine-grained detail, enabling expressive and natural-sounding speech synthesis. 2) Since continuous features inherently capture variations in tone, pitch, and emphasis, they are well-suited for prosody control and emotional TTS. 3) Continuous representations are robust to information loss and can avoid quantization artifacts, allowing reconstruction of smooth audio. GAN-based [115, 143, 144] and diffusion-based methods [149, 148] often utilize continuous feature representations, i.e., mel-spectrograms. However, continuous representations are typically computationally demanding and require large-scale models and training datasets, especially in high-resolution audio synthesis.
Discrete Tokens. In discrete token-based TTS, the intermediate acoustic features (e.g., quantized units or phoneme-like tokens) are discrete values, similar to words or phonemes in languages. These are often produced using quantization techniques or learned token embeddings, such as HuBERT [166] and SoundStream [170]. The advantages of discrete tokens are: 1) Discrete tokens can encode phonemes or sub-word units, making them concise and computationally efficient to handle. 2) Discrete tokens often allow TTS systems to require fewer samples to learn and generalize, as compared with continuous representations, since the representations are compact and simplified. 3) Using discrete tokens simplifies cross-modal TTS applications like voice cloning or text prompt-based TTS, as they map well to text-like representations such as LLM tokens. LLM-based [78, 103, 106, 105] and zero-shot TTS methods [17, 78, 87] often adopt discrete tokens as their acoustic features. However, discrete representation learning may result in information loss or lack the nuanced details that can be captured in continuous representations.
Speech Quantization vs. Tokenization. It is worth noting that quantization and tokenization serve distinct purposes in speech processing. Quantization is primarily used for high-fidelity compression, reducing the precision of numerical representations (e.g., from 32-bit floating point to 8-bit integers) while preserving model performance. In speech synthesis, quantization is often used in waveform generation (e.g., codec-based approaches like EnCodec [171]) and neural vocoders to compress audio signals without significant loss of perceptual quality. Tokenization, on the other hand, is a discretization process that segments continuous data into meaningful units. In speech tasks, tokenization extracts semantically relevant representations such as phonemes, characters, or learned speech units (e.g., HuBERT [166] and Wav2Vec 2.0 [165]). This makes tokenization particularly suitable for speech-to-text (ASR), TTS, and multimodal NLP tasks, where aligning speech with textual information is crucial. Tokenization also facilitates training language models on speech data by enabling linguistic or learned unit-based processing rather than raw audio waveform modeling. Table II summarizes popular open-source speech quantization and tokenization methods. Tables IV and III summarize the types of acoustic features of representative methods.
@@ III Uncontrollable TTS
The development of uncontrollable text-to-speech (UC-TTS) systems represents a significant shift from traditional, linguistics-based synthesis to modern, data-driven deep learning techniques. This shift highlights the integration of both local and global information to produce speech with human-like quality and naturalness. This section briefly investigates UC-TTS methods, emphasizing the role of local and global information in enhancing speech fidelity and expressiveness.
In the context of UC-TTS, the term “uncontrollable” refers to the absence of explicit control mechanisms for speech features such as emotion, timbre, and speaking style. Despite this, the goal is to achieve natural, fluid speech while mitigating issues like mispronunciations and omissions.
@@ III-A Early Approaches: Statistical Models
Early TTS systems relied on statistical models such as HMMs [64, 65] and early neural network-based parametric methods [110, 111]. These models operated at the frame level, using acoustic models and vocoders for text-to-speech conversion. Notable contributions from Tokuda et al. [177] employed HMMs for statistical parametric synthesis, focusing on local features like phonemes, accents, and prosody to improve speech naturalness.
@@ III-B Sequence-to-Sequence Models
The emergence of sequence-to-sequence models represents a significant breakthrough by removing the need for explicit linguistic features, thereby enabling the capture of the nuances of human speech. Models such as Tacotron [74] and Tacotron 2 [75] utilize RNNs with attention mechanisms to effectively model the complex, nonlinear nature of speech sequences. These innovations enable the tuning of speech parameters, enhancing prosody and rhythm by modeling entire utterances instead of isolated phonetic units.
Building on these advancements, Deep Voice 3[121] introduces a fully convolutional sequence-to-sequence architecture that significantly accelerates training speed compared to RNN-based models. This approach achieves training times an order of magnitude faster, enabling scalability to handle large datasets. Additionally, the position-augmented attention mechanism in Deep Voice 3 enhances the naturalness of synthesized speech, achieving competitive mean opinion scores, especially when paired with advanced neural vocoders like WaveNet. This development not only improves training efficiency but also enhances the scalability and naturalness of TTS systems.
@@ III-C Transformer-based Models
III-D Integrating Flow and Diffusion Models
Recent advancements have shifted toward integrating global information within end-to-end architectures to enhance speech naturalness and coherence. Flow-based models like Glow-TTS[132] and Flow-TTS[131] exemplify this by employing invertible transformations that maintain the balance between local precision and global coherence. These architectures enable the synthesis of high-fidelity speech by modeling complex dependencies across the entire utterance, thus improving the overall fluidity and naturalness of the generated speech.
Moreover, the introduction of diffusion models in TTS, such as WaveGrad 2[178], highlights the shift toward models that can iteratively refine speech output. These models use score-matching and diffusion processes to generate speech directly from phoneme sequences, effectively capturing both local nuances and overarching global patterns. The iterative nature of these models allows for adjustments that enhance the quality of the synthesized audio, accommodating variations in speech without explicit control over specific attributes.
The integration of adversarial training and VAEs further exemplifies the evolution toward incorporating global information. Systems like VITS[159] leverage these techniques to enhance expressiveness and naturalness by learning complex mappings between text and speech. This approach allows the model to manage variations in prosody and rhythm that are inherently derived from the textual input, aligning with the objectives of UC-TTS to produce diverse and natural speech.
@@ IV Controllable TTS
In this section, we first review recent TTS work from the perspective of model architecture, followed by a detailed discussion of control strategies in controllable TTS, which is the core part of this survey. Current model architectures can be broadly classified into two main categories: The first is the non-autoregressive (NAR) generative models, which are based on HMMs, neural networks, VAEs, diffusion models, flow matching, and other NAR techniques. The second category relies on autoregressive (AR) codec language models, which typically quantize speech into discrete tokens and use decoder-only models to autoregressively generate these tokens. We summarize the NAR-based and AR-based controllable TTS methods in Table III and Table IV, respectively.
@@
FastSpeech[15] Transformer WaveGlow [179] MelS 2019.05
FastSpeech 2[180] Transformer Parallel WaveGAN [143] MelS 2020.06
@@
Parallel Tacotron[181] Transformer + CNN WaveRNN [136] MelS 2020.10
StyleTagging-TTS [182] Transformer + CNN HiFi-GAN [115] MelS 2021.04
SC-GlowTTS [183] Transformer + Flow HiFi-GAN MelS 2021.06
Meta-StyleSpeech [184] Transformer MelGAN [144] MelS 2021.06
DelightfulTTS [185] Transformer + CNN HiFiNet [185] MelS 2021.11
YourTTS [82] Transformer + Flow HiFi-GAN LinS 2021.12
StyleTTS [88] CNN + RNN HiFi-GAN MelS 2022.05
GenerSpeech [90] Transformer + Flow HiFi-GAN MelS 2022.05
Cauliflow[186] BERT + Flow UP WaveNet [187] MelS 2022.06
CLONE[188] Transformer + CNN WaveNet [73] MelS + LinS 2022.07
PromptTTS [101] BERT + Transformer HiFi-GAN MelS 2022.11